# Demystifying Behavior-Based Malware Detection at Endpoints

### v1

@@ Appendix C Case Studies on Sandbox-Specific Artifacts
Appendix D Fine-Tuning a Sandbox-Based Model on Endpoint Traces
@@
Appendix E More Details on the Leaderboard
The following artifacts will be released to the participants:
Sandbox Dataset. We will release the training portion of our sandbox datasets in Table 3, stored in the standardized format discussed in Section 4.1. We avoid releasing the testing portion to prevent participants from obtaining an impractical advantage by training their detectors on it (will be released at a later date). This will level the playing field for participants and ensure they all have access to the same sandbox data for development. Further, as this data contains traces from two sandboxes, participants can leverage our observation in Section 5 and tune their models on a second sandbox, not seen during training, for potentially higher endpoint performance. Moreover, they can also apply invariance learning techniques, which have been shown promising in Section 7.4.
Sample Metadata. We will release the metadata relating to the training samples, including their SHA-256 hashes, ground truth labels, family tags (if malware), and first-seen timestamps. Moreover, our metadata also tags the source of a sample, e.g., EMBER (anderson2018ember, ), SOREL (harang2020sorel20m, ), or our endpoint dataset. This allows participants to easily obtain realistic priors over malware families for endpoint detection, and make use of our findings and improvements in Section 6.3. Moroever, they can also create realistic testing distributions over samples using these malware family priors (and avoid the problems we discussed in Section 6.
Appendix F Additional Tables and Figures
@@
Figure 6. Statistics on the endpoint traces in our dataset.
@@
Table 13. Top malware families (top) and benign (bottom) publishers in our EP (left) and SB (right) test sets, along with their shares in each dataset.
Soft Labeling Function Refer to caption
Figure 7. Our function for assigning probabilistic (soft) labels to each sample based on the number of VirusTotal detections.
@@
Table 14. For Wannacry traces in our SB1, SB2, and EP datasets, we first measure the ratio of traces with at least one and no IOC. We then measure the average prediction score of our NGR models in SB1->EP, SB2->EP and EP->EP on these traces.
Rank Test Set Trained on SB1 Trained on SB2 Trained on EP SB1->SB1 SB1->SB2 SB1->EP SB2->SB1 SB2->SB2 SB2->EP EP->EP NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT SB1 99.1 99.0 98.9 93.8 91.3 92.6 76.1 70.7 72.8 89.5 83.4 89.1 97.7 94.1 96.9 72.0 68.2 70.8 85.6 85.2 84.3 SB2 99.0 98.6 98.8 94.9 93.2 93.5 76.6 71.3 74.4 81.7 71.5 86.4 98.7 98.5 98.2 71.6 66.7 72.7 85.4 85.9 84.8 EP 99.0 98.6 98.7 92.9 91.2 92.9 78.4 74.1 74.9 84.4 71.8 88.3 98.0 97.8 98.0 74.6 72.0 73.4 87.5 86.8 86.8
Table 15. The performance (AUC%) of three ML approaches (NGR, HYB, ATT) in seven detection scenarios (based on Table 2). In each row, the models are ranked based on the test set, and the average AUC of top-$20$ models is reported.
Trace Length Bias of the Sandbox-Based Model
Figure 8. Comparing the correlations between trace length and malware-ness prediction scores of the model. Malware ratio is the ground truth ratio of malware traces of a certain length among all traces of that length in a dataset.
Figure 9. Prediction score standard deviations on the endpoint traces of each sample in the EP test set; for a model in SB1->EP (left), and EP->EP (right).
Fine-Tuning a Sandbox-Based Model on Endpoint Traces Refer to caption
Figure 10. The results of fine-tuning a sandbox-based model in SB->EP on endpoint traces. The dashed line indicates our best model SB->EP. Experiments on NGR.

### v2

@@ Appendix E The Impact of Sandbox Evasion
Appendix F Fine-Tuning a Sandbox-Based Model on Endpoint Traces
@@
Figure 6: The results of fine-tuning a sandbox-based model in SB->EP on endpoint traces. The dashed line indicates our best model SB->EP. Experiments on NGR.
Sel. Test Set Trained on SB1 Trained on SB2 Trained on EP SB1->SB1 SB1->SB2 SB1->EP SB2->SB1 SB2->SB2 SB2->EP EP->EP NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT NGR HYB ATT SB1 99.1 99.0 98.9 93.8 91.3 92.6 76.1 70.7 72.8 89.5 83.4 89.1 97.7 94.1 96.9 72.0 68.2 70.8 85.6 85.2 84.3 SB2 99.0 98.6 98.8 94.9 93.2 93.5 76.6 71.3 74.4 81.7 71.5 86.4 98.7 98.5 98.2 71.6 66.7 72.7 85.4 85.9 84.8 EP 99.0 98.6 98.7 92.9 91.2 92.9 78.4 74.1 74.9 84.4 71.8 88.3 98.0 97.8 98.0 74.6 72.0 73.4 87.5 86.8 86.8
TABLE XIII: The performance (AUC%) of three ML approaches (NGR, HYB, ATT) in seven detection scenarios (based on Table II). The models are selected using the test set in each row, and the average AUC of the top 20 models is reported.
Appendix G More Details on the Artifact Release
We will release the following artifacts to the community (visit our website for more details https://malwaredetectioninthewild.github.io/):
Sandbox Dataset. We will release the training portion of our sandbox datasets in Table III, stored in the standardized format discussed in Section IV-A. We avoid releasing the testing portion publicly to prevent researchers who wish to participate in our realistic evaluation leaderboard from obtaining an impractical advantage by training their detectors on it (will be released if requested in exchange for being excluded from the leaderboard). This will level the playing field for participants and ensure they all have access to the same sandbox data for development. Further, as this data contains traces from two sandboxes, participants can leverage our observations in Section V and tune their hyper-parameters on a second sandbox, not seen during training, for improving the endpoint performance or train their models on data from both sandboxes. They can also apply invariant learning techniques, which have shown promising results in Section VII-D.
Sample Metadata. We will release the metadata relating to the training samples in all datasets, including their SHA-256 hashes, ground truth labels, family tags (if malware), and first-seen timestamps. Additionally, our metadata also annotates the source of a sample, e.g., EMBER [49], SOREL [53], or our endpoint dataset. This allows researchers to compute realistic priors over malware families for endpoint detection and make use of our findings and improvements in Section VI-C. Moreover, they can also create realistic testing distributions over samples using these malware family priors (and avoid the problems we discussed in Section VI.
Appendix H Additional Tables and Figures
@@
Figure 7: Statistics on the endpoint traces in our dataset.
Soft Labeling Function Refer to caption
Figure 8: Our function for assigning probabilistic (soft) labels to each sample based on the number of VirusTotal detections.
Figure 9: Prediction score standard deviations on the endpoint traces of each sample in the EP test set; for a model in SB1->EP (left), and EP->EP (right).
@@
TABLE XIV: Top malware families (top) and benign (bottom) publishers in our EP (left) and SB (right) test sets, along with their shares in each dataset.
@@
TABLE XV: For Wannacry traces in our SB1, SB2, and EP datasets, we first measure the ratio of traces with at least one and no IOC. We then measure the average prediction score of our NGR models in SB1->EP, SB2->EP and EP->EP on these traces.